OpenStudy (anonymous):

Solve: 2r^2-11x+5=0 Thanks :) p.s. explain?

OpenStudy (anonymous):

Do you know the quadratic formula?

hero (hero):

The variables should be the same, right?

OpenStudy (anonymous):

Agreed.

OpenStudy (anonymous):

Good eye.

hero (hero):

Good eye? It was blatantly obvious, lol :P

OpenStudy (anonymous):

oh right! My teacher made a typo, lol. they should both be r's. and no, I don't remember anything from algebra 1! Which is bad :/

hero (hero):

How do you know they shouldn't both be x's?

OpenStudy (anonymous):

well, letters don't really matter. whichever you replace, they would still be the same problem. just different letters.

OpenStudy (anonymous):

I think he's being humorous, sakigirl.

OpenStudy (anonymous):

can someone just help me? lol. i'm kinda stuck on this question >_<

hero (hero):

2r^2-11r+5=0 2r^2-10r-1r+5 = 0 2r(r-5)-1(r-5) = 0 (r-5)(2r-1) = 0 r = 1/2 r = 5

OpenStudy (anonymous):

Hero is right.

OpenStudy (mimi_x3):

You can also use the quadratic formula, which is faster i guess. \[\large x=\frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2(a)} \]

hero (hero):

Nope, I'm faster. The quad formula is whack.

OpenStudy (anonymous):

Hahaha!

hero (hero):

Mimi uses it because she can't factor :P

OpenStudy (mimi_x3):

Nope, well..the using the quadratic you won't need to think much as to factor where you mgiht make stupid mistakes.

hero (hero):

Mimi, you forgot to put the line that protects your discriminant :P

OpenStudy (mimi_x3):

what line?

hero (hero):

The square root line.....(DUH) :P

OpenStudy (mimi_x3):

wtf..there's nothing wrong :/ You're confusing me :/

hero (hero):

lol, just teasing. Don't cry on me :P

OpenStudy (anonymous):

HERO. one last question. i beg.

OpenStudy (mimi_x3):

Oh..gawd, confusing a poor kid that just woke up :'(

hero (hero):

Here we go again with the "poor little kid" retreat line you always pull :P

OpenStudy (mimi_x3):

You dont have the right to alternate my phrases i said "poor kid" NOT "poor little kid" :P

hero (hero):

poor kid = poor little kid To me it's the exact same thing. A sympathy plea

OpenStudy (anonymous):

kwell, I'm just gonna ask. 2c^2+c-10=0

hero (hero):

Mimi can use her quad form to help you

OpenStudy (mimi_x3):

Excuse me, its not a "sympathy plea" i never told you to be pity me. GOsh people these days who dont have common sense.

OpenStudy (anonymous):

i don't remember crap about the quad form hah.

OpenStudy (mimi_x3):

Because Hero is unable to factor it :P

OpenStudy (mimi_x3):

\[\large c=\frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2(a)} \] \[\large c=\frac{-(1)\pm\sqrt{(1)^{2}-4(2)(-10)}}{2(2)} \]

hero (hero):

2c^2+c-10=0 2c^2+5c-4c-10 = 0 c(2c+5)-2(2c+5) = 0 (2c+5)(c-2) = 0 c = -5/2, 2

hero (hero):

Mimi jinxed me.

OpenStudy (mimi_x3):

lol, i told you that you will make stupid mistakes :P The factor method is such a FAIL!

OpenStudy (anonymous):

Mimi made careless mistakes! :O

hero (hero):

Quad Form is an even bigger fail, especially when you depend on it so much Mimi :P

hero (hero):

I know about 5 different factor methods

hero (hero):

None of which include the dreaded quad form

OpenStudy (anonymous):

how would you use the quad form on questions like 25m^2-64=0??

OpenStudy (mimi_x3):

What? You dont have the right to say that its a fail since i got it right in the first go, not like you who go it WRONG! EVen during exams it would be a waste of time to do it again :P

OpenStudy (anonymous):

a.k.a binomials.

hero (hero):

25m^2 - 64 = 0 25m^2 = 64 m^2 = 64/25 m = sqrt{64/25} m = plus or minus 8/5

OpenStudy (mimi_x3):

Btw, quadratic formula can only be used for ax^3+bx+c

OpenStudy (anonymous):

Oh right, I totally knew that.

OpenStudy (mimi_x3):

ax^2*

hero (hero):

Yeah, the quad form is LIMITED

hero (hero):

While I was working hard to provide your solution sakigirl. Mimi, started making useless comments

OpenStudy (anonymous):

hahah, forgetting the line of the square root. Oh silly her.

OpenStudy (mimi_x3):

I was trying to help her know a better method to solve it which is quicker. I was being nice

hero (hero):

I was being nicer :P

OpenStudy (mimi_x3):

whaat? forgettng the line of the square root?

hero (hero):

lol

OpenStudy (anonymous):

Yeah! and Hero commented on it

hero (hero):

lol

OpenStudy (mimi_x3):

lol, sorry to say i think you should scroll up, i never make those carelless mistakes.

hero (hero):

you just made one, then auto corrected it. Putting ax^3 , then ax^2. Such silly mistakes :P

OpenStudy (mimi_x3):

its called a TYPO, you make stupid mistakes as well argh..i should never comment in questions that you are in anymore.